Hydrogen cyanide can be prepared by reacting of methane, CH4, with ammonia by the following reaction.

CH4(g) + NH3(g) →HCN(g) +3H2(g)

What is the heat of reaction at constant pressure?  Express your answer in kJ

Useful information:

N2(g) +3H2(g) → 2NH3(g); ΔH = -91.8 kJ

C(graphite) +2H2(g) → CH4(g); ΔH = -74.9 kJ

H2(g) +2C(graphite) +N2(g) → 2HCN(g); ΔH = 270.3 kJ
Answer #1

Lets number the reaction as 0, 1, 2, 3 from top to bottom

required reaction should be written in terms of other reaction

This is Hess Law

required reaction can be written as:

reaction 0 = -0.5 * (reaction 1) -1 * (reaction 2) +0.5 * (reaction 3)

So, ΔHo rxn for required reaction will be:

ΔHo rxn = -0.5 * ΔHo rxn(reaction 1) -1 * ΔHo rxn(reaction 2) +0.5 * ΔHo rxn(reaction 3)

= -0.5 * (-91.8) -1 * (-74.9) +0.5 * (270.3)

= 255.95 KJ

Answer: 255.95 KJ